Bob's age B and Tom's age T are related by the expression =2T-11. When Bob is 35 years old, how old is Tom?
Answer by ankor@dixie-net.com(22740)

Bob's age B and Tom's age T are related by the expression = 2T-11.
2T - 11 = B
When Bob is 35 years old, how old is Tom?
2T - 11 = 35
2T = 35 + 11
2T = 46
T = 46/2
T = 23 yrs is Tom's age
:
:
Check:
B = 2(23) - 11
B = 46 - 11
B = 35